OpenStudy (anonymous):

find an equation for f^-1(x), the inverse function verify that your equation is correct by showing that f(f^-1(x))=x and f^-1(f(x))=x f(x)=x+3 f(x)=4x

OpenStudy (anonymous):

\[f(x)=x+3 == y = x+3\] \[y=x+3\] \[x=y-3\] the inverse is: \[f(x)^{-1}=x-3\] \[f(x)=4x\] \[y=4x\] \[x=y/4\] the inverse is: \[f(x)^{-1}=x/4\] remember, for a function to have inverse it has to be one to one function, a one to one function must pass vertical line test, that is, a each x has unique y

OpenStudy (anonymous):

\[f(x-3)=(x-3)+3=x\] \[f(x/4)=4(x/4)=x\]

OpenStudy (anonymous):